Circular No. 7034

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA SCORPII 1998 W. Liller, Viña del Mar, Chile, reports his discovery of an apparent

nova (mag 6.9) on Technical Pan photographs taken with PROBLICOM on Oct. 21.0266, 21.0293, and 21.0389 UT with an 85-mm camera lens (+ orange filter); he gave the location of the new star as α = 17^{h}55^{m}.4, δ = −31^{o}00′ (equinox 2000.0). Nothing brighter than mag 11.5 apppears at this location on photographs taken on Oct. 17. G. Garradd, Loomberah, N.S.W., reports the following precise position and magnitude from an unfiltered CCD image obtained on Oct. 22.409 with a 0.45-m f/5.4 reflector: α = 17^{h}55^{m}24^{s}.99, δ = −31^{o}01′41′′.5, 7.4. A. Jones, Nelson, New Zealand, reports visual magnitude 8.4 on Oct. 22.43. No star within 15′′ of the new star is present in the USNO A1.0 or Guide Star catalogues.

SUPERNOVA 1998eh IN ESO 74-G9 J. Maza, Department of Astronomy, University of Chile, reports the

discovery by M. Wischnjewsky, on a T-Max 400 film taken by L. E. González with the Maksutov telescope at Cerro El Roble on Oct. 15.078 UT, of a supernova (B ∼ 18) located at α = 20^{h}48^{m}57^{s}.82, δ = −69^{o}05′05′′.7 (equinox 2000.0), which is 5′′.2 east and 25′′.5 north of the nucleus of the spiral galaxy ESO 74-G9. The object was confirmed via B and V direct images taken by D. González on Oct. 18.06 using the YALO telescope at Cerro Tololo. SN 1998eh is not visible on a film obtained with the Cerro El Roble telescope on Sept. 28.025 (B > 19.5). The object is 0.5 mag brighter on Oct. 18 (compared to Oct. 15).

SUPERNOVA 1998ei IN ANONYMOUS GALAXY Maza also reports the discovery by Roberto Antezana, taken as above

by L. E. González on Oct. 15.201 UT, of a supernova (B ∼ 18.5) located at α = 23^{h}32^{m}03^{s}.75, δ = −39^{o}33′18′′.4 (equinox 2000.0), which is 11′′.8 west and 10′′.9 south of the nucleus of the host spiral galaxy. The object was confirmed via B and V direct images taken by D. González on Oct. 17 using the YALO telescope at Cerro Tololo. SN 1998ei is not visible on a film obtained with the Cerro El Roble telescope on Sept. 28.223 (B > 19.5).

COMET P/1998 S1 Visual m_{1} estimates by K. Hornoch, Lelekovice, Czech Republic (0.35-

m reflector): Oct. 20.89 UT, 14.3; 21.82, 14.0.

1998 October 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green